Q-IV Calculate the percentage energy transmitted at water to steel interface, given that the acoustic impedance for steel is 46.7, and 1.48 for water. Q-V What is the smallest defect you could detect with 2 MHz probe inspecting a steel specimen with a velocity of 6x10 cm /sec
Answer #1

(a)

Amount of energy transmitted = total amount of energy - amount of energy reflected

The amount of energy reflected = ( (Z1 - Z2) ÷ (Z1+Z2))2

Z1 = acoustic impedance of steel = 46.7

Z2= acoustic impedance of air. = 1.48

Therefore ,

The reflected energy = ( ( 46.7 - 1.48) ÷ (46.7+1.48))2

.   = ( 45.22 ÷ 48.18)2

= 0.88

Percentage of energy reflected = 0.88 × 100

= 88%

Total amount of energy = 100%

The amount of transmitted energy = 100- 88

= 12%

(b)